OpenStudy (anonymous):

How long would it take for 1.50 mol of water at 100.0 ∘C to be converted completely into steam if heat were added at a constant rate of 23.0J/s ?

OpenStudy (aaronq):

\(\sf rate=\dfrac{heat}{time}\) \(\sf time=\dfrac{heat}{rate}=\dfrac{n*H_{vap}}{rate}=\dfrac{1.5~mol*40680~J/mol}{23.0~J/s}=2653.04~s\)